# AceCoder: Acing Coder RL via Automated Test-Case Synthesis

### v2

@@ 1 Introduction
Therefore, we curate AceCode-89K, on which we trained our reward models: AceCode-RM-7B and AceCode-RM-32B. Comprehensive experiments of best-of-N sampling show that AceCode-RM can significantly boost existing LLM’s performance on coding benchmarks. For example, AceCode-RM-7B can improve the performance of Llama-3.1-8B-Instruct by an average of $8.4$ points across the 4 coding benchmarks, i.e. HumanEval (Liu et al., 2023), MBPP (Liu et al., 2023), BigCodeBench (Zhuo et al., 2024) and LiveCodeBench (Jain et al., 2024). Even for the stronger coder model Qwen2.5-Coder-7B-Instruct, our "7B+7B" combination still gets an average of $2.6$ improvements. AceCode-RM-32B is even more powerful, which pushes the former two numbers to $10.7$ and $4.7$ respectively, showcasing the effectiveness of AceCode-RM.
@@
To our knowledge, this is the first work to perform reward model training and reinforcement learning for code generation using a fully automated pipeline that synthesizes large-scale reliable tests. We believe our AceCode-89K will unlock the potential of RL training for code generation models and help the community to further push the boundaries of LLM’s coding abilities.
@@ 2 Methodology
In this section, we will introduce the overall methodology of AceCoder. We begin with formulations of the problems we are investigating, including reward model training and reinforcement learning for LLMs. We then elaborate on how we synthesize the test cases and construct the AceCode-89K. Finally, we explain how we perform the reinforcement learning using our AceCode-RM trained on the AceCode-89K.
@@ 2.1 Problem Formulation
3 AceCode-89K
To be able to train a reward model specifically designed for code generation, the first thing is to synthesize reliable test cases for each coding problem and use them as training signals. In this section, we explain the whole procedure of constructing AceCode-89K step by step. We show the overall statistics in Table 1.
@@ 3 AceCode-89K
We start from existing coding datasets with provided question $\mathbf{x}$ and corresponding program $\mathbf{y}$. Specifically, we combine Magicoder-Evol-Instruct111ise-uiuc/Magicoder-Evol-Instruct-110K, Magicoder-OSS-Instruct-75K222ise-uiuc/Magicoder-OSS-Instruct-75K, and StackPyFunction333bigcode/stack-dedup-python-fns as our seed dataset. We only keep the questions written in Python that contain either a function or a class, resulting in a total of 124K entries. We find that these datasets contain highly noisy questions that could not be easily evaluated using test cases. Therefore, we feed every question-solution pair ($\mathbf{x}$, $\mathbf{y}$) into a GPT-4o-mini (Hurst et al., 2024) to propose a refined LeetCode-style question $\mathbf{x_{r}}$ with highly structured instructions. Meanwhile, we also prompt it to ‘imagine’ around 20 test cases $(t_{1},...,t_{m})$ for each refined coding question $\mathbf{x_{r}}$ based on its understanding of the expected behavior of the desired program. See prompt template used in subsection A.1. Please note that we do not use the program solution $\mathbf{y}$ from the existing datasets at all in our final curated AceCode-89K. These datasets are purely used as seeds to help LLM formulate well-structured coding problems.
@@
These ‘imagined’ test cases generated from the LLM contain severe hallucinations. To filter out those hallucinated test cases, we facilitated a stronger coder model Qwen2.5-Coder-32B-Instruct (Hui et al., 2024a) as a proxy to perform quality control. Specifically, we prompt it for each $\mathbf{x_{r}}$ to generate a program $\mathbf{y}^{\prime}$ and then run these programs over the test cases to approximate their quality. We removed all test cases $t_{i}$ where the generated solution program $\mathbf{y}^{\prime}$ could not pass. Furthermore, we removed questions with fewer than 5 tests after filtering, as these questions might be overly ambiguous. With the above filtering, we constructed the AceCode-89K with 89.4K distinct coding questions and 1.39M cleaned test cases, as represented by $(\mathbf{x_{r}},(t_{1},...,t_{m_{c}}))$, where $m_{c}$ represents the number of test cases after filtering.
@@
# Examples 27,853 26,346 35,223 89,422
# Avg Test Cases 14.77 16.11 15.79 15.56
@@
Table 1: Dataset statistics of AceCode-89K before and after test-case filtering.
@@
We propose to use the Bradley-Terry model to train the reward model as defined in Equation 1. Therefore, we need to construct (question, [positive program, negative program]) data from AceCode-89K. Specifically, we sample programs $(\mathbf{y}^{1},...,\mathbf{y}^{n})$ from existing models (e.g. Llama-3.1 (Grattafiori et al., 2024)) w.r.t $\mathbf{x}_{r}$ and utilize the test-case pass rate to distinguish positive and negative programs. Since the pass rate $s_{i}$ for the sampled program $\mathbf{y}^{i}$ can be any number between $[0,1]$, a minor difference in pass rate may not represent that one program is more accurate than another. Therefore, instead of using $\mathbbm{1}[s_{i}>s_{j}]$ to select the preference pairs, we have thus modified the selection rules to be:
@@ 4.1 Reward Model Training Setup
We mainly use Qwen2.5-Coder-7B-Instruct  [^4] Qwen/Qwen2.5-Coder-7B-Instruct as the backbone of the reward model and sample 16 responses from it for each question in AceCode-89K. Finally, following the rule defined in  Equation 3, around 300K preference pairs were created out of 46,618 distinct questions (37.34% of the total questions) that have at least one pair satisfying the condition, and other questions are not used.
@@
DeepSeek-V2.5 90.2 83.5 87.6 74.1 53.2 29.1 48.9 27.0 41.8 59.5
@@ 4.2 Reinforcement Learning Setup
We perform RL training from three policy models: Qwen2.5-7B-Instruct  [^5] Qwen/Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Base  [^6] Qwen/Qwen2.5-Coder-7B and Qwen2.5-Coder-7B-Instruct. Two types of reward can be used, i.e. the trained reward model AceCode-RM-7B and the rule-based reward, i.e. pass rate over the test cases in AceCode-89K. During training, we set the pass rate to be a binary reward, which is 1.0 when all test cases passed, otherwise 0. This is similar to the verfiable reward used in Tulu3 (Lambert et al., 2024a) and DeepSeek-R1 (Guo et al., 2025). Similar to DeepSeek-R1 (Guo et al., 2025), we also experiment with RL from the base model because SFT may cause the search space of the model to be stuck in the local minimum. Since coding is also a highly verifiable task like math, we include the Qwen2.5-Coder-7B-Base in our experiments.
We have trained different policy model backbones with different rewards, resulting in 6 RL models in total. All the RL-tuning are based on OpenRLHF (Hu et al., 2024). We adopt the Reinforcement++ (Hu, 2025) algorithm instead of PPO to improve the training efficiency without training the value model. It’s also proved to be more stable than PPO and GRPO. We train our model on a subsampled hard version of AceCode-89K, where we keep the 25% of the questions with lower average pass rates and higher variance. This is to ensure the question is hard and that the sampled programs are diverse enough. For the training hyperparameters, we set the rollout batch size to 256, and 8 programs are sampled from per question. The training batch size is 128 with a learning rate of 5e-7. All the models are trained for 1 episode and finished in 6 hours on 8 x H100 GPUs.
@@ 4.4 Main Results
Here we show the reward model evaluation results through Best-of-N.
Best-of-N Results
@@
When starting from the Qwen2.5-Coder-Instruct-7B itself, we can still observe improvements, especially when using the rule-based reward. For example, we get more than $3.4$ improvement on BigCodeBench-Full-Hard. Using the reward model for RL can also bring $3.2$ improvement on MBPP, making it ($86.0$) only 1.6 points behind compared to DeepSeek-V2.5 ($87.6$). This highlights the charm of self-improvement given the reward model backbone is the same with the initial policy model.
@@ 5.4 Reinforcement Learning for LLM
Since the introduction of Reinforcement Learning from Human Feedback (RLHF)(Ouyang et al., 2022b), it has been extensively applied to enhance LLM capabilities in tasks such as conversational interactions and mathematical reasoning(Yang et al., 2024b). Popular reinforcement learning algorithms, including PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and Reinforcement++ (Hu, 2025), have been used to fine-tune models according to reward signals generated either by reward models (Shao et al., 2024) or pre-defined rule-based rewards (Guo et al., 2025; Wang et al., 2025). Despite the fact that coding is inherently a verifiable task, reinforcement learning has seen limited application in code generation due to challenges in defining meaningful and scalable reward signals. The most relevant prior work, CodeRL (Le et al., 2022), leverages the APPS dataset (Hendrycks et al., 2021) to generate rewards based on test case evaluations. However, APPS contains only 5,000 examples, with most problems having just a single test case, making it insufficient for scalable RL-based training. Our work is the first to automate high-quality test case synthesis in code domain.

### v3

@@ 1 Introduction
Therefore, we curate AceCode-87K, on which we trained our reward models: AceCode-RM-7B and AceCode-RM-32B. Comprehensive experiments of best-of-N sampling show that AceCode-RM can significantly boost existing LLM’s performance on coding benchmarks. For example, AceCode-RM-7B can improve the performance of Llama-3.1-8B-Instruct by an average of $8.4$ points across the 4 coding benchmarks, i.e. HumanEval (Liu et al., 2023), MBPP (Liu et al., 2023), BigCodeBench (Zhuo et al., 2024) and LiveCodeBench (Jain et al., 2024). Even for the stronger coder model Qwen2.5-Coder-7B-Instruct, our "7B+7B" combination still gets an average of $2.6$ improvements. AceCode-RM-32B is even more powerful, which pushes the former two numbers to $10.7$ and $4.7$ respectively, showcasing the effectiveness of AceCode-RM.
@@
To our knowledge, this is the first work to perform reward model training and reinforcement learning for code generation using a fully automated pipeline that synthesizes large-scale reliable tests. We believe our AceCode-87K will unlock the potential of RL training for code generation models and help the community to further push the boundaries of LLM’s coding abilities.
@@ 2 Methodology
In this section, we will introduce the overall methodology of AceCoder. We begin with formulations of the problems we are investigating, including reward model training and reinforcement learning for LLMs. We then elaborate on how we synthesize the test cases and construct the AceCode-87K. Finally, we explain how we perform the reinforcement learning using our AceCode-RM trained on the AceCode-87K.
@@ 2.1 Problem Formulation
3 AceCode-87K
To be able to train a reward model specifically designed for code generation, the first thing is to synthesize reliable test cases for each coding problem and use them as training signals. In this section, we explain the whole procedure of constructing AceCode-87K step by step. We show the overall statistics in Table 1.
@@ 3 AceCode-87K
We start from existing coding datasets with provided question $\mathbf{x}$ and corresponding program $\mathbf{y}$. Specifically, we combine Magicoder-Evol-Instruct111ise-uiuc/Magicoder-Evol-Instruct-110K, Magicoder-OSS-Instruct-75K222ise-uiuc/Magicoder-OSS-Instruct-75K, and StackPyFunction333bigcode/stack-dedup-python-fns as our seed dataset. We only keep the questions written in Python that contain either a function or a class, resulting in a total of 124K entries. We find that these datasets contain highly noisy questions that could not be easily evaluated using test cases. Therefore, we feed every question-solution pair ($\mathbf{x}$, $\mathbf{y}$) into a GPT-4o-mini (Hurst et al., 2024) to propose a refined LeetCode-style question $\mathbf{x_{r}}$ with highly structured instructions. Meanwhile, we also prompt it to ‘imagine’ around 20 test cases $(t_{1},...,t_{m})$ for each refined coding question $\mathbf{x_{r}}$ based on its understanding of the expected behavior of the desired program. See prompt template used in subsection A.1. Please note that we do not use the program solution $\mathbf{y}$ from the existing datasets at all in our final curated AceCode-87K. These datasets are purely used as seeds to help LLM formulate well-structured coding problems.
@@
These ‘imagined’ test cases generated from the LLM contain severe hallucinations. To filter out those hallucinated test cases, we facilitated a stronger coder model Qwen2.5-Coder-32B-Instruct (Hui et al., 2024a) as a proxy to perform quality control. Specifically, we prompt it for each $\mathbf{x_{r}}$ to generate a program $\mathbf{y}^{\prime}$ and then run these programs over the test cases to approximate their quality. We removed all test cases $t_{i}$ where the generated solution program $\mathbf{y}^{\prime}$ could not pass. Furthermore, we removed questions with fewer than 5 tests after filtering, as these questions might be overly ambiguous. With the above filtering, we constructed the AceCode-87K with 87.1K distinct coding questions and 1.38M cleaned test cases, as represented by $(\mathbf{x_{r}},(t_{1},...,t_{m_{c}}))$, where $m_{c}$ represents the number of test cases after filtering.
@@
# Examples 26,920 25,862 34,367 87,149
# Avg Test Cases 15.14 16.33 16.08 15.87
@@
Table 1: Dataset statistics of AceCode-87K before and after test-case filtering.
@@
We propose to use the Bradley-Terry model to train the reward model as defined in Equation 1. Therefore, we need to construct (question, [positive program, negative program]) data from AceCode-87K. Specifically, we sample programs $(\mathbf{y}^{1},...,\mathbf{y}^{n})$ from existing models (e.g. Llama-3.1 (Grattafiori et al., 2024)) w.r.t $\mathbf{x}_{r}$ and utilize the test-case pass rate to distinguish positive and negative programs. Since the pass rate $s_{i}$ for the sampled program $\mathbf{y}^{i}$ can be any number between $[0,1]$, a minor difference in pass rate may not represent that one program is more accurate than another. Therefore, instead of using $\mathbbm{1}[s_{i}>s_{j}]$ to select the preference pairs, we have thus modified the selection rules to be:
@@ 4.1 Reward Model Training Setup
We mainly use Qwen2.5-Coder-7B-Instruct  [^4] Qwen/Qwen2.5-Coder-7B-Instruct as the backbone of the reward model and sample 16 responses from it for each question in AceCode-87K. Finally, following the rule defined in  Equation 3, around 300K preference pairs were created out of 46,618 distinct questions (37.34% of the total questions) that have at least one pair satisfying the condition, and other questions are not used.
@@
RLEF-8B - 67.5 - 57.0 - - - - - -
RLEF-70B - 78.5 - 67.6 - - - - - -
PPOCoder-7B 78.7 - 67.0 - - - - - - -
StepCoder-7B 76.8 - 63.8 - - - - - - -
CodeGemma-7B 60.5 - 55.2 - - - - - - -
DSTC-33B 79.9 72.0 82.5 70.4 51.6 22.3 41.0 18.2 - -
@@ 4.2 Reinforcement Learning Setup
We perform RL training from three policy models: Qwen2.5-7B-Instruct  [^5] Qwen/Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Base  [^6] Qwen/Qwen2.5-Coder-7B and Qwen2.5-Coder-7B-Instruct. Two types of reward can be used, i.e. the trained reward model AceCode-RM-7B and the rule-based reward, i.e. pass rate over the test cases in AceCode-87K. During training, we set the pass rate to be a binary reward, which is 1.0 when all test cases passed, otherwise 0. This is similar to the verfiable reward used in Tulu3 (Lambert et al., 2024a) and DeepSeek-R1 (Guo et al., 2025). Similar to DeepSeek-R1 (Guo et al., 2025), we also experiment with RL from the base model because SFT may cause the search space of the model to be stuck in the local minimum. Since coding is also a highly verifiable task like math, we include the Qwen2.5-Coder-7B-Base in our experiments.
We have trained different policy model backbones with different rewards, resulting in 6 RL models in total. All the RL-tuning are based on OpenRLHF (Hu et al., 2024). We adopt the Reinforcement++ (Hu, 2025) algorithm instead of PPO to improve the training efficiency without training the value model. It’s also proved to be more stable than PPO and GRPO. We train our model on a subsampled hard version of AceCode-87K, where we keep the 25% of the questions with lower average pass rates and higher variance. This is to ensure the question is hard and that the sampled programs are diverse enough. For the training hyperparameters, we set the rollout batch size to 256, and 8 programs are sampled from per question. The training batch size is 128 with a learning rate of 5e-7. All the models are trained for 1 episode and finished in 6 hours on 8 x H100 GPUs.
@@ 4.4 Main Results
Here we show the experimental results of reward model and RL-trained model.
RM Results
@@
When starting from the Qwen2.5-Coder-Instruct-7B itself, we can still observe improvements, especially when using the rule-based reward. For example, we get more than $3.4$ improvement on BigCodeBench-Full-Hard. Using the reward model for RL can also bring $3.2$ improvement on MBPP. This highlights the charm of self-improvement given the reward model backbone is the same with the initial policy model. We compare our method with other RL-based models like RLEF (Chen et al., 2024), PPOCoder (Shojaee et al., 2023a), StepCoder (Dou et al., 2024b), DSTC (Liu et al., 2024c), etc. We show that our 7B model is able to beat these competitors across the evaluation benchmarks.
@@ 5.4 Reinforcement Learning for LLM
Reinforcement Learning from Human Feedback (RLHF)(Ouyang et al., 2022b) has been widely adopted to enhance the capabilities of large language models (LLMs) in various tasks, including conversational interactions and mathematical reasoning(Yang et al., 2024b). Reinforcement learning (RL) algorithms such as PPO(Schulman et al., 2017), GRPO(Shao et al., 2024), and Reinforcement++(Hu, 2025) have been employed to fine-tune models using reward signals derived from either learned reward models(Shao et al., 2024) or predefined rule-based heuristics (Guo et al., 2025; Wang et al., 2025).
Given that coding is an inherently verifiable task, recent studies have explored RL techniques that leverage direct execution accuracy as a reward signal. PPOCoder (Shojaee et al., 2023b) and CodeRL (Le et al., 2022) demonstrated the effectiveness of PPO-based RL for coding tasks, while RLEF (Gehring et al., 2024) extended this approach to multi-turn settings by incorporating execution feedback at each step. StepCoder (Dou et al., 2024a) refined the reward mechanism by assigning rewards at a more granular level, considering only successfully executed lines of code. Additionally, DSTC (Liu et al., 2024d) explored the application of Direct Preference Optimization (DPO) to code generation by using self-generated test cases and programs.
Despite these advancements, most prior RL-based approaches for coding have been constrained by the use of pre-annotated datasets such as APPS (Hendrycks et al., 2021), which consists of only 5,000 examples, with most problems having a single test case. This limited data availability poses challenges to scalable RL training. Furthermore, the potential of reward models for coding remains largely unexplored. In this work, we address these limitations by automatically synthesizing test cases and leveraging trained reward models for reinforcement learning, demonstrating the scalability and effectiveness of our approach.